Help please
20 of 40 (0 complete This Question: 2 pts A survey of 200 people was conducted to see if they smoke. The results are shown below Female 74 17 91 Tota 161 39 200 87 Doesnt smoke 22 Total 109 If a person is selected at random, what is the probability that the person is a male or is a smoker? OA. 1785 O B. 0423 O C. correct answer is not given D. 0.915 it a person is selected at random, what is the probablity that the person is a male given that the person was a smoker? A. 0.798 B comect answer is not given С. 0435 D. 0.540 Cick to select your answer
Answer #1

Solution:

P(male or smoker) =P(Male)+P(Smoker)-P(male and smoker)

=( 109+161-87)/200 = 183/200 = 0.915

So it's answer is D. I.e. 0.915

Solution()

P(male/smoker) = P(male and smoker)/p(smoker) = 87/200/161/200 = 87/161 = 0.54

So it's answer is D. I.e. 0.540